# Feature-Based Interpretable Optimization

### v1

@@ Abstract
For optimization models to be used in practice, it is crucial that users trust the results. A key factor in this aspect is the interpretability of the solution process. A previous framework for inherently interpretable optimization models used decision trees to map instances to solutions of the underlying optimization model. Based on this work, we investigate how we can use more general optimization rules to further increase interpretability and at the same time give more freedom to the decision maker. The proposed rules do not map to a concrete solution but to a set of solutions characterized by common features. To find such optimization rules, we present an exact methodology using mixed-integer programming formulations as well as heuristics. We also outline the challenges and opportunities that these methods present. In particular, we demonstrate the improvement in solution quality that our approach offers compared to existing frameworks for interpretable optimization and we discuss the relationship between interpretability and performance. These findings are supported by experiments using both synthetic and real-world data.
Keywords: interpretable optimization; interpretability and explainability; data-driven optimization; contextual optimization; optimization under uncertainty
@@ 1.1 Motivation
The widespread availability of easy-to-use, off-the-shelf machine learning tools has dramatically expanded their use. Yet users often perceive these tools as black-boxes. This is because while the optimization model and the solution processes may be understandable, it is difficult to explain how the solution came about and how it might change in response to different parameters. Consequently, the importance of interpretability and explainability in computer-aided decision processes has increased significantly in recent years, especially in areas where large datasets form the basis. In optimization, however, only recently have researchers started to aim for more interpretable or explainable approaches. One reason for this delay might be that, despite the availability of solvers, expert knowledge is still crucial for appropriately modeling optimization problems. These experts possess a deep understanding of both the modeling and solution techniques involved. Consequently, there has been little intrinsic motivation for mathematicians and optimization specialists to prioritize more interpretable or explainable solution techniques.
@@
However, this perspective is not shared by all stakeholders (i.e., those affected by the decision) who must accept and implement the optimization results. For instance, workers who are directly affected by optimized decisions, such as those scheduled or guided by the outcome of an optimization process. They need to understand these decisions so as to help them prepare for future outcomes, e.g.potential overtime or subsequent planning decisions. The lack of transparent decision processes or explanations can lead to discontent and loss of trust, resulting in poor adoption of optimized decisions and ultimately rendering the optimization process ineffective. Interpretable decisions also allow the users to adapt to changing circumstances on the fly, leading to better outcomes.
@@
Therefore, to encourage actions based on optimization programs and build trust in their outcomes, clear, practical, and easily understandable explanations are essential. Justifying why a system recommends certain actions is more important to stakeholders than explaining how the solution is obtained [MB00, YJ95]. Hence, providing the underlying optimization model may not be as useful, since “disclosure alone does not guarantee that anyone is paying attention or is able to accurately interpret the information; more complex information is more likely to be unexamined or misunderstood” [Pra05].
@@
Moreover, offering adequate information about the company, its decisions, and the worker’s role within it can enhance their identification with the company [SPVR01]. To this end, effective and transparent communication is crucial in fostering trust and positive relationships between employees and the organization [Raw08, MS14, YMF19, WXW18]. Conversely, reliance on complex and hard-to-understand decision models can foster skepticism and a reluctance to adopt decision-support systems, even if these models have been shown to improve decision-making performance [ACC+06, KDBL+09].
@@
Therefore, it is essential for managers and decision-makers to provide insights into decisions derived from optimization processes. As a first step, incorporating workers into the feedback cycle of the modeling process is crucial, as providing explanations during this process enhances the feedback loop [CSGK19]. However, even after a model is agreed upon, stakeholders still need clear and easily understandable explanations of when and why different solutions are considered. This is evident in decision support systems, where decision-makers or stakeholders must explain outcomes, such as in loan approval processes [SYX+20, SMF21] or medical consultations [SCDS77, Swa85, LSG+19]. Ultimately, the European “right to explanation” ensures that users can request an explanation for any algorithmic decision made about them [GF17].
@@
With this mindset, our goal is to provide straightforward, easily comprehensible rules that make it possible to deduce a result from the instance itself. In other words, we are aiming for an interpretable optimization method. This differs from explaining a solution, which merely requires a post hoc justification of the result. It is important to note that any interpretable approach is inherently explainable, as one can simply reference the rule used to derive the solution. However, the reverse is rarely true: explainable approaches usually are not interpretable [Rud19].
@@
In this paper, we rethink the framework presented in [GH23]. There, the authors aimed to find comprehensible optimization rules that map cost scenarios to solutions. However, this approach has the drawback of not allowing the user to deviate from the provided solution, making it difficult to adapt to unexpected situations. Furthermore, when describing a solution comprehensibly in practice, one typically focuses on the most relevant features rather than detailing every minor aspect. Hence, in this paper, we extend the interpretable optimization framework to make use of features. In particular, any given vector of instance features is mapped (in a comprehensible way) to a vector of solution features. We call the corresponding set of solutions that exhibit these features a meta-solution.
@@ 1.2 Related Literature
Interpretability and explainability have only recently gained attention in the field of optimization. In contrast, the broader field of artificial intelligence (AI) has seen a rapid increase in interest regarding explainable AI [AB18] and more interpretable approaches [Rud19] over the last decade. This shift in AI underscores the importance of transparency and trust in complex systems, a principle that is equally vital in optimization.
@@
To provide context, we first discuss what constitutes “easily comprehensible” descriptions, acknowledging the subjectivity of this terminology, before we present literature aimed at enhancing comprehensibility. Several studies have investigated which representation types are most comprehensible for human users [Mil19, Arz21]. These studies suggest that decision trees and classification rules [BSH+10, HDM+11, Fre14], as well as linear models [SL97, PSLB15, UR16], are particularly understandable, with sparsity being a significant factor contributing to their comprehensibility [Gle16]. To achieve comprehensible mappings from instances to solutions as well as to balance quality and interpretability, several approaches have been proposed. These include post hoc simplification to improve interpretability [NZJT12], enforcing linear representations [BHSR15], and semantically constraining the generated rules [Hun16]. Overall, most approaches claim interpretability based on the reduced size of the resulting rules and expressions.
Interpretable optimization.
@@
Interpretability calls for an easily comprehensible mapping from instances to solutions. In this sense, linear programming can be considered an interpretable framework: given an instance, one can apply the simplex algorithm (an “easily comprehensible” rule) to find the solution. However, as previously discussed, the ability to comprehend such methods greatly depends on the individual’s domain knowledge. Therefore, more generally comprehensible mappings are needed to achieve broader interpretability. This necessitates optimization rules that balance quality and interpretability. Consequently, the goal of obtaining interpretable optimization rules can be viewed as finding easily comprehensible heuristics. This approach is closely linked to the field of generation hyper-heuristics [DKÖB20]. Early approaches in this field already aimed to ensure interpretability by creating representations with limited size [BHKW07, NZ17]. Various strategies have been developed to harmonize quality and interpretability, such as simplifying heuristics post hoc [NZJT12], using linear representations for clarity [BHSR15], and imposing semantic constraints on the rules [Hun16]. Recent applications of hyper-heuristics that also strive for interpretability include routing policies [WMZ19], dispatching rules [FFA22], and online combinatorial optimization problems [ZBQ+22]. Note that most of these approaches to generate heuristics use genetic programming [BHK+19] or other search algorithms, which stands in contrast to the optimization approach presented in this paper.
@@
Furthermore, approaches tailored for specific settings have been introduced, such as a method for stopping in stochastic systems [CM22] and a mixed integer programming approach to obtain interpretable tree-based models for clustering [BOW21]. However, these methods only assign class labels (stop/continue, cluster membership) in an interpretable manner and do not apply to general optimization settings.
@@
In contrast, the framework presented in this paper offers an interpretable optimization rule that can be applied to any optimization problem. It allows for easily comprehensible mappings from instances to understandable solution spaces described by features. This approach generalizes the method presented in [GH23], which had the limitation of only being able to map to concrete solutions, thus reducing flexibility when encountering new scenarios.
@@
In contrast to interpretability, explainability involves post hoc justifications for why a particular solution was chosen, without necessarily providing true insights into the underlying mechanism. In optimization, necessary and sufficient optimality conditions and sensitivity analysis can serve as forms of explanation. Additionally, contrastive explanations can help clarify why the found solution is optimal. These explanations often require a mathematical background and a deep understanding of the optimization model.
To address this complexity, various methods for enhancing explainability in optimization have emerged in recent years, drawing inspiration from AI techniques. In [DBCDC+24] the authors outline how AI methods solving operations research problems should be used responsibly, also considering explainability as one critical aspect. To advance explainable optimization, initially, the focus was domain-specific, such as argumentation-based explanations for planning problems [CMP19, ODV20] and scheduling problems [ČLMT19, ČLL21]. Later, more general methods were developed for dynamic programming [EK21], multi-stage stochastic optimization [TBBN22], and linear optimization [KBH24]. Furthermore, data-driven explainability frameworks were introduced in [AGH+24] and [FPV23]. Additionally, for the case of multi-objective optimization problems, where interaction with a decision maker is crucial to balance different criteria and reach the best compromise on the Pareto frontier, several methods have been proposed to enhance explainability [SSG18, MALM22, CGMS24, Mis24].
@@
Optimization under uncertainty is highly relevant to our work, as we aim for an optimization rule that has to work well for any (unknown) data that we might observe in the future. In particular, problems with a two-stage setting are closely related. In our setting, the first stage consists of finding an appropriate optimization rule, such that for the anticipated scenarios, the realized solution in the second stage (that has to adhere to the optimization rule) minimizes the overall decision criteria, which can be for example the worst-case or the Laplace criteria. A very close connection can be drawn to so-called decision rules [GKW19], where in a multistage setting, decisions that need to be taken in later stages are merged into a decision rule, which has to be determined in the first decision stage. In particular, a decision rule specifies the reaction to a scenario and can therefore be viewed as an explanation for how a scenario affects a decision. However, the primary motivation for using decision rules is not to enhance the comprehensibility of decisions in later stages, but to obtain approximate decisions in a reasonable time. The selection of function classes for decision rules is based mainly on computational performance and approximation quality. Also in the realm of optimization under uncertainty, a connection can be drawn to $K$-adaptability [BK17, BK18], where the best $K$ solutions are selected in the first decision stage, with the decision of which solution to choose in a specific scenario being deferred until the scenario is revealed. Notably, there is no rule provided for mapping scenarios to these solutions and furthermore, these solutions are static, in contrast to our setting of meta-solutions. Only in recent studies has a connection been explored between $K$-adaptability and decision rules [SGW20, VGY20], but these efforts have prioritized improving computational properties rather than enhancing comprehensibility.
@@
Furthermore, predictive prescriptions [BK20] and in particular prescriptive trees [BDM19, JAGV21] were introduced, prescribing decisions based on observation. Prescriptive trees are closely related to decision trees and hence their interpretability is frequently emphasized. The main difference to our approach is that our candidates assigned to the leaves result from an optimization process rather than from classification. In particular, we do not assign prefixed solutions but select suitable ones via optimization.
@@
Another related research field is the smart “predict then optimize” framework [EG22], which links the prediction of parameter data to the optimization process for which this data is intended. Instead of minimizing the loss of the predicted parameters, the focus is on the loss in cost that results from using predicted parameters compared to true parameters. There have also been advancements in making the predict-and-optimize framework more explainable [BCGV23]. Our approach on the other hand merges prediction and optimization into a single mechanism, mapping features directly to solutions rather than to parameters that are then used to find an optimal solution. In particular, we do not predict parameters at all; instead, we propose (meta-)solutions.

### v3

# Feature-Based Interpretable Surrogates for Optimization

@@ Abstract
For optimization models to be used in practice, it is crucial that users trust the results. A key factor in this aspect is the interpretability of the solution process. A previous framework for inherently interpretable optimization models used decision trees to map instances to solutions of the underlying optimization model. Based on this work, we investigate how we can use more general optimization rules to further increase interpretability and, at the same time, give more freedom to the decision-maker. The proposed rules do not map to a concrete solution but to a set of solutions characterized by common features. To find such optimization rules, we present an exact methodology using mixed-integer programming formulations as well as heuristics. We also outline the challenges and opportunities that these methods present. In particular, we demonstrate the improvement in solution quality that our approach offers compared to existing interpretable surrogates for optimization, and we discuss the relationship between interpretability and performance. These findings are supported by experiments using both synthetic and real-world data.
Keywords: interpretable surrogates; interpretability and explainability; data-driven optimization; contextual optimization; optimization under uncertainty
@@ 1.1 Motivation
The widespread availability of easy-to-use, off-the-shelf machine learning tools has dramatically expanded their use. Yet users often perceive these tools as black boxes. Consequently, the importance of interpretability and explainability has increased significantly in recent years, especially in areas where large datasets form the basis. For optimization problems, however, only recently have researchers started to aim for more interpretable or explainable approaches. One reason for this delay might be that, despite the availability of solvers, expert knowledge is still crucial for appropriately modeling optimization problems. These experts possess a deep understanding of both the modeling and solution techniques involved. Consequently, there has been little intrinsic motivation for mathematicians and optimization specialists to prioritize more interpretable or explainable solution techniques.
@@
However, this perspective is not shared by all stakeholders especially those who must accept and implement the optimization results. For instance, workers who are directly affected by optimized decisions, such as those scheduled or guided by the outcome of an optimization process. They need to understand these decisions so as to help them prepare for future outcomes, e.g., potential overtime or subsequent planning decisions. The lack of transparent decision processes or explanations can lead to discontent and loss of trust, resulting in poor adoption of optimized decisions and ultimately rendering the optimization process ineffective. Interpretable solution approaches also allow users to adapt to changing circumstances on the fly, leading to better outcomes.
@@
Therefore, to encourage acceptance of actions based on optimization programs and build trust in their outcomes, clear, practical, and easily understandable explanations are essential. Justifying why a system recommends certain actions is more important to stakeholders than explaining how the solution is obtained (Mao and Benbasat, 2000; Ye and Johnson, 1995). Hence, providing the underlying optimization model may not be as useful since “disclosure alone does not guarantee that anyone is paying attention or is able to accurately interpret the information; more complex information is more likely to be unexamined or misunderstood” (Prat, 2005).
@@
Moreover, offering adequate information about the company, its decisions, and the worker’s role within it can enhance their identification with the company (Smidts et al., 2001). To this end, effective and transparent communication is crucial in fostering trust and positive relationships between employees and the organization (Rawlins, 2008; Men and Stacks, 2014; Yue et al., 2019; Wang et al., 2018). Conversely, reliance on complex and hard-to-understand decision models can foster skepticism and a reluctance to adopt decision-support systems, even if these models have been shown to improve decision-making performance (Arnold et al., 2006; Kayande et al., 2009).
@@
Therefore, it is essential for managers and decision-makers to provide insights into decisions derived from optimization processes. As a first step, incorporating workers into the feedback cycle of the modeling process is crucial, as providing explanations during this process enhances the feedback loop (Chakraborti et al., 2019). Already in this phase, interpretable surrogates unlock several practical benefits beyond transparency: they make model faults (e.g., missing constraints, miscalibrated thresholds, unintended biases) immediately apparent to non-experts, simplifying debugging, and they enable straightforward counterfactual explanations by showing which minimal feature changes would alter a recommendation. However, even after a model is agreed upon, stakeholders still need clear and easily understandable explanations of when and why different solutions are considered. This is evident in decision support systems, where decision-makers or stakeholders must explain outcomes, such as in loan approval processes (Sachan et al., 2020; Strich et al., 2021) or medical consultations (Scott et al., 1977; Swartout, 1985; Lamy et al., 2019). Ultimately, the European “right to explanation” ensures that users can request an explanation for any algorithmic decision made about them (Goodman and Flaxman, 2017).
The main contribution of this paper is not only to generate interpretable surrogates for optimization, but to explore how their inputs and outputs should be presented. We argue that, in practice, conveying only high-level solution features—rather than a fully specified solution—better aligns with human decision‐making. For instance, in the shortest-path setting, route-choice studies show that travelers rely on a handful of landmarks to guide their routes, rendering strict optimal paths poor predictors of actual behavior (Manley et al., 2015b, a). This insight suggests that feature-based recommendations support intuitive decision-making and that offering suboptimal yet easily comprehensible guidance aligns with real-world practice, since humans rely on heuristics and seldom pursue strictly optimal solutions. Moreover, by prescribing a set of essential features (e.g., “pass through landmarks A, B, and C”, “avoid city center”), we leverage users’ domain knowledge to fill in the implementation details, reducing the need for micromanagement. Such flexibility also accommodates minor environmental changes: the features can flexibly map to slightly different realizations without compromising the outlined solution scope. In contrast, prescribing a single rigid solution risks being overly constraining—potentially leading users away from near-optimal outcomes—whereas solution features strike a balance between guidance and adaptability. Similarly, representing a new instance using features is often more practical in real-world settings. Beyond this convenience, observable features—such as seasonality or weather—can significantly influence decisions, and omitting them can lead to systematically biased or inconsistent outcomes (Ban and Rudin, 2019).
@@
With this mindset, our goal is to provide straightforward, easily comprehensible, feature-based rules that make it possible to deduce a result from the instance itself. In other words, we are aiming for an interpretable surrogate for the optimization process that indicates meaningful solution features while empowering users to determine the precise implementation. This differs from explaining a solution, which merely requires a post hoc justification of the result. It is important to note that any interpretable approach is inherently explainable, as one can simply reference the rule used to derive the solution. However, the reverse is rarely true: explainable approaches usually are not interpretable (Rudin, 2019).
@@
In this paper, we rethink the framework presented in Goerigk and Hartisch (2023). There, the authors aimed to find comprehensible optimization rules that map cost scenarios to solutions. However, this approach has the drawback of not allowing the user to deviate from the provided solution, making it difficult to adapt to unexpected situations. Furthermore, when describing an instance or a solution comprehensibly in practice, one typically focuses on the most relevant features rather than detailing every minor aspect. Hence, in this paper, we extend the interpretable optimization framework to make use of features. In particular, any given vector of instance features is mapped (in a comprehensible way) to a vector of solution features. We call the corresponding set of solutions that exhibit these features a meta-solution. Figure 1 provides a schematic comparison of our approach and the related procedures used in data-driven optimization.
Figure 1: Comparison of the proposed feature-based framework with related approaches for solving optimization problems. In our method, the traditional use of a black-box solver is replaced by an interpretable surrogate model, which is first trained on available data. For a new problem instance, the surrogate specifies a restricted solution space using comprehensible features, enabling the user to select a solution suitable for their situation.
@@ 1.2 Related Literature
Interpretability and explainability have only recently gained attention in the field of optimization. In contrast, the broader field of artificial intelligence (AI) has seen a rapid increase in interest regarding explainable AI (Adadi and Berrada, 2018) and more interpretable approaches (Rudin, 2019) over the last decade. This shift in AI underscores the importance of transparency and trust in complex systems, a principle that is equally vital in optimization.
@@
To provide context, we first discuss what constitutes “easily comprehensible” descriptions, acknowledging the subjectivity of this terminology, before we present literature aimed at enhancing comprehensibility. Several studies have investigated which representation types are most comprehensible for human users (Miller, 2019; Arzamasov, 2021). These studies suggest that decision trees and classification rules (Baehrens et al., 2010; Huysmans et al., 2011; Freitas, 2014), as well as linear models (Setiono and Liu, 1997; Parker et al., 2015; Ustun and Rudin, 2016), are particularly understandable, with sparsity being a significant factor contributing to their comprehensibility (Gleicher, 2016). To achieve comprehensible mappings from instances to solutions as well as to balance quality and interpretability, several approaches have been proposed. These include post hoc simplification to improve interpretability (Nguyen et al., 2012), enforcing linear representations (Branke et al., 2015), and semantically constraining the generated rules (Hunt, 2016). Overall, most approaches claim interpretability based on the reduced size of the resulting rules and expressions.
Interpretable optimization approaches.
@@
Interpretability calls for an easily comprehensible mapping from instances to solutions. In this sense, linear programming can be considered an interpretable framework: given an instance, one can apply the simplex algorithm (an “easily comprehensible” rule) to find the solution. However, as previously discussed, the ability to comprehend such methods greatly depends on the individual’s domain knowledge. Therefore, more generally comprehensible mappings are needed to achieve broader interpretability. This necessitates optimization rules that balance quality and interpretability. Consequently, the goal of obtaining interpretable optimization rules can be viewed as finding easily comprehensible heuristics. This approach is closely linked to the field of generation hyper-heuristics (Drake et al., 2020). Early approaches in this field already aimed to ensure interpretability by creating representations with limited size (Burke et al., 2007; Nguyen and Zhang, 2017). Various strategies have been developed to harmonize quality and interpretability, such as simplifying heuristics post hoc (Nguyen et al., 2012), using linear representations for clarity (Branke et al., 2015), and imposing semantic constraints on the rules (Hunt, 2016). Recent applications of hyper-heuristics that also strive for interpretability include routing policies (Wang et al., 2019), dispatching rules (Ferreira et al., 2022), and online combinatorial optimization problems (Zhang et al., 2022). Note that most of these approaches to generate heuristics use genetic programming (Burke et al., 2019) or other search algorithms, which stands in contrast to the optimization approach presented in this paper.
@@
Furthermore, approaches tailored for specific settings have been introduced, such as a method for stopping in stochastic systems (Ciocan and Mišić, 2022) and a mixed integer programming approach to obtain interpretable tree-based models for clustering (Bertsimas et al., 2021). However, these methods only assign class labels (stop/continue, cluster membership) in an interpretable manner and do not apply to general optimization settings.
@@
In contrast, the framework proposed in this paper provides an interpretable optimization rule that serves as a surrogate for the optimization process and can be applied to any optimization problem. It allows for easily comprehensible mappings from instances to understandable solution spaces described by features. This approach generalizes the method presented in Goerigk and Hartisch (2023), which had the limitation of only being able to map to concrete solutions, thus reducing flexibility when encountering new scenarios. In the rest of this work, we will also refer to this framework as the solution-based approach and to optimal decision rules generated using it as solution-based interpretable surrogates. Other works, which pursue more general concepts, deal with the generation of decision trees which map to solution strategies (Bertsimas and Stellato, 2021) and the robustification of the solution-based approach (Goerigk et al., 2024).
@@
In contrast to interpretability, explainability involves post hoc justifications for why a particular solution was chosen without necessarily providing true insights into the underlying mechanism. In optimization, necessary and sufficient optimality conditions and sensitivity analysis can serve as forms of explanation. Additionally, contrastive explanations can help clarify why the found solution is optimal. These explanations often require a mathematical background and a deep understanding of the optimization model.
To address this complexity, various methods for enhancing explainability in optimization have emerged in recent years, drawing inspiration from AI techniques. In De Bock et al. (2024), the authors outline how AI methods solving operations research problems should be used responsibly, also considering explainability as one critical aspect. To advance explainable optimization, several domain-specific approaches evolved, such as argumentation-based explanations for planning problems (Collins et al., 2019; Oren et al., 2020), scheduling problems (Čyras et al., 2019, 2021), and workforce allocation (Povéda et al., 2024). Later, more general methods were developed for dynamic programming (Erwig and Kumar, 2021), multi-stage stochastic optimization (Tierney et al., 2022), and linear optimization (Kurtz et al., 2024). Furthermore, data-driven explainability frameworks were introduced in Aigner et al. (2024) and Forel et al. (2023), where the importance of selecting suitable features is a crucial aspect (Aigner et al., 2025). Additionally, for the case of multi-objective optimization problems, where interaction with a decision maker is crucial to balance different criteria and reach the best compromise on the Pareto frontier, several methods have been proposed to enhance explainability (Sukkerd et al., 2018; Misitano et al., 2022; Corrente et al., 2024; Misitano, 2024).
@@
Optimization under uncertainty is highly relevant to our work, as we aim for an optimization rule that has to work well for any (unknown) data that we might observe in the future. In particular, problems with a two-stage setting are closely related. In our setting, the first stage consists of finding an appropriate optimization rule, such that for the anticipated scenarios, the realized solution in the second stage (that has to adhere to the optimization rule) minimizes the overall decision criteria, which can be for example, the worst-case or the Laplace criteria. A very close connection can be drawn to so-called decision rules (Bertsimas and Georghiou, 2015; Georghiou et al., 2019), where in a multistage setting, decisions that need to be taken in later stages are merged into a decision rule, which has to be determined in the first decision stage. In particular, a decision rule specifies the reaction to a scenario and can, therefore be viewed as an explanation for how a scenario affects a decision. However, the primary motivation for using decision rules is not to enhance the comprehensibility of decisions in later stages but to obtain approximate decisions in a reasonable time. The selection of function classes for decision rules is based mainly on computational performance and approximation quality. Also, in the realm of optimization under uncertainty, a connection can be drawn to $K$-adaptability (Buchheim and Kurtz, 2017; Hanasusanto et al., 2015; Malaguti et al., 2022), where the best $K$ solutions are selected in the first decision stage, with the decision of which solution to choose in a specific scenario being deferred until the scenario is revealed. Notably, there is no rule provided for mapping scenarios to these solutions, and furthermore, these solutions are static, in contrast to our setting of meta-solutions. Only in recent studies has a connection been explored between $K$-adaptability and decision rules (Subramanyam et al., 2020; Vayanos et al., 2020), but these efforts have prioritized improving computational properties rather than enhancing comprehensibility.
@@
Furthermore, predictive prescriptions (Bertsimas and Kallus, 2020) and, in particular, prescriptive trees (Bertsimas et al., 2019; Jo et al., 2021) were introduced, prescribing decisions based on observation. Prescriptive trees are closely related to decision trees, and hence their interpretability is frequently emphasized. The main difference to our approach is that our candidates assigned to the leaves result from an optimization process rather than from classification. In particular, we do not assign prefixed solutions but select suitable ones via optimization.
@@
Another related research field is contextual optimization (Sadana et al., 2025), in particular the smart “predict then optimize” framework (Elmachtoub and Grigas, 2022), which links the prediction of parameter data to the optimization process for which this data is intended. Instead of minimizing the loss of the predicted parameters, the focus is on the loss in cost that results from using predicted parameters compared to true parameters. There have also been advancements in making the predict-and-optimize framework more explainable (Blanquero et al., 2023). Furthermore, the integrated use of decision trees was investigated to increase the interpretability of the approach (Elmachtoub et al., 2020). Our approach, on the other hand, merges prediction and optimization into a single mechanism, mapping features directly to solutions rather than to parameters that are then used to find an optimal solution. In particular, we do not predict parameters at all; instead, we propose (meta-)solutions.